# Entropy-Aware Structural Alignment for Zero-Shot Handwritten Chinese Character Recognition

### v1

@@ 4.1.2 Implementation Details
The proposed framework is implemented using PyTorch on a workstation equipped with four NVIDIA RTX 4090 GPUs. For the visual stream, we utilize a randomly initialized ResNet-34 [8] backbone, which is trained from scratch on the training set to learn domain-specific handwritten features.
The semantic radical encoder is fine-tuned using weights from the CLIP (ViT-B/32) [22] image encoder. This fine-tuning process is conducted prior to the main backbone training, following the protocol detailed in our preliminary work [20]. Consequently, all radical features are extracted via this pre-trained CLIP encoder to construct a static feature dictionary. During the main network training, the CLIP parameters are frozen and not updated, serving solely for efficient vector retrieval.
The entire model is optimized end-to-end using the Adadelta optimizer with a batch size of 32. We employ a MultiStepLR scheduler, where the learning rate is initialized to 0.1 and decayed to 0.01 and 0.001 at the 15th and 20th epochs, respectively. Regarding key hyperparameters, the neighbor count $K$ for the Top-K Semantic Feature Fusion strategy is set to 5 based on our ablation studies (see Section 4.3.3).
@@ 4.2.1 Performance on Zero-shot Character Recognition
The core contribution of this work is the ability to recognize unseen characters by leveraging radical-based semantic priors. We evaluate our method against a comprehensive list of SOTA approaches. To ensure a thorough comparison, selected baselines cover diverse structural modeling strategies, including:
  • 1.
    Radical-based methods: ranging from attention-based aggregation (HDE [3], SLD [4]) to recent hierarchical tree models (SideNet [16], HierCode [44]).
  • 2.
    Stroke-based methods: which utilize stroke-level decomposition (STAR [41], RSST [39]).
  • 3.
    Multi-modal frameworks: which integrate visual-semantic alignment or auxiliary modalities (CCR-CLIP [40], UCR-RRM [15]).
@@
Table 1 summarizes the Top-1 accuracy on the ICDAR 2013 target set (Unseen) under the zero-shot setting, where the number of seen training classes $m$ varies from 500 to 2,755. Our Entropy-Aware Structural Alignment Network establishes a new SOTA, consistently outperforming all existing methods across different data regimes.
Superiority over Advanced Structural Models: Compared to the latest hierarchical methods such as HierCode and SideNet, which also employ radical trees, our method achieves a remarkable performance leap of +11.75% and +17.66% respectively on the full dataset ($m=2755$). This significant gap highlights that structure alone is insufficient; the introduction of the Information Entropy Prior is the decisive factor that enables the model to distinguish fine-grained unseen characters by suppressing low-information structural noise.
Advancement over Linear and Stroke Baselines: Compared to our preliminary conference version CCR-CLIP, which relies on linear radical sequences, the proposed dual-view tree extension yields a solid +5.37% improvement, validating the necessity of topological modeling. Furthermore, our method significantly outperforms stroke-based approaches like RSST by over 20%, suggesting that for zero-shot semantic transfer, radical-level representation provides a more robust semantic anchor than stroke-level features.
@@ 4.2.2 Generalization to Full-Set Recognition
While our primary focus is zero-shot learning, we also evaluate the proposed framework on the standard fully supervised setting (using the complete CASIA-HWDB 1.0/1.1 training set) to verify its general applicability. Table 2 compares our method with SOTA fully supervised approaches. The comparison considers two key dimensions: the type of Auxiliary Information (e.g., Glyphs, Radicals, Strokes) and the Similarity Metric Strategy.
@@
Performance vs. Efficiency Analysis: As shown in Table 2, our method achieves a highly competitive accuracy of 97.30%.
  • 1.
@@
Impact of Metric Strategy: The results also highlight the importance of the matching mechanism. Most existing methods (e.g., HDE, SLD) employ Cosine Similarity, which is computationally efficient but often fails to capture complex non-linear cross-modal dependencies. By introducing Cross-Modal Attention followed by Top-K Aggregation, our framework can better align visual features with semantic embeddings, leading to more robust recognition than fixed-distance metrics.
@@ 4.3 Ablation Studies
4.3.1 Effectiveness of Entropy-Aware Position Embedding
To isolate the contribution of the proposed Entropy-Aware Position Embedding (EAPE), we conduct a progressive ablation study comparing it against standard positional encoding schemes on the zero-shot radical recognition task. The baseline employs a simple one-hot radical encoding without any structural position information. We then incrementally incorporate the radical tree topology and various position embedding strategies to measure their specific impacts.
The quantitative results are summarized in Table 4. As observed, the introduction of the Radical Tree topology alone (without explicit position indices) triggers the most significant performance leap, boosting the accuracy from 16.53% to 23.23% at $m=500$. This confirms that the hierarchical connection itself provides essential structural context that a flat baseline lacks. However, simply adding standard geometric encodings—such as Sinusoidal PE or RoPE—to this tree structure yields only marginal gains (e.g., Sinusoidal PE improves by merely $\sim$0.4% over the tree-only baseline). This saturation suggests that treating character structure purely as a geometric sequence is suboptimal, as it fails to account for the varying semantic importance of different nodes.
In contrast, our Entropy-Aware mechanism effectively breaks this bottleneck, achieving the highest accuracy of 24.54% ($m=500$) and 55.04% ($m=1500$). By utilizing multiplicative interaction to modulate the positional embeddings, our method amplifies the feature magnitude of high-entropy radicals while dampening ubiquitous particles. This empirical evidence validates that the ”information value” of a radical is a more critical attention prior than its geometric coordinate for recognizing unseen characters.
@@ 4.4 Analysis of Few-Shot Adaptability and Data Efficiency
While zero-shot recognition demonstrates the model’s ability to reason from semantic definitions, real-world applications often permit a ”cold start” with a minimal number of support samples. To evaluate the data efficiency of our framework, we investigate the performance gains when providing $K$ reference samples per unseen category, specifically setting $K\in\{1,2,5\}$.
Rapid Adaptation Performance: The performance trends and detailed numerical results are presented in Figure 7. The results exhibit a compelling ”Rapid Adaptation” characteristic. Most notably, providing a single visual prototype ($K=1$) triggers a massive performance leap. For $m=500$, the accuracy surges from 24.54% (zero-shot) to an impressive 92.41%. This implies that a single reference image is sufficient to bridge the vast majority of the domain gap. Increasing the support set to $K=2$ further refines the decision boundary, pushing the accuracy to 96.62%. However, as $K$ increases to 5, the performance saturates at 97.93%, showing diminishing returns. This saturation curve confirms that our model is highly sample-efficient: it captures the core character identity with as few as two examples, eliminating the need for large-scale data fitting typical of traditional deep learning models.
Mechanism of Efficiency: We attribute this superior efficiency to the Entropy-Aware Structural Prior. In conventional few-shot learning, the model must implicitly infer ”what is discriminative” by statistically comparing multiple samples, a process that inherently requires a larger $K$ to suppress noise. In contrast, our framework operates with a pre-computed ”structural blueprint.” The information entropy explicitly highlights high-value radicals (e.g., specific strokes distinguishing similar characters) before any visual samples are seen. Consequently, when the first support image ($K=1$) arrives, the attention mechanism immediately ”locks onto” these critical regions, creating a shortcut in the optimization landscape. Effectively, the radical tree serves as a robust coordinate system, allowing the model to align distorted handwritten features with canonical structures instantly. This transforms the learning process from ”statistical exploration” to ”structural registration,” enabling robust recognition with minimal visual evidence.
@@ 4.5 Visualization of Cross-Modal Reasoning Process
To provide deeper insights into the internal mechanisms of the proposed Entropy-Aware Structural Alignment Network, we visualize the inference process for a challenging unseen sample, opening the “black box” of cross-modal reasoning. We select a handwritten image of the character “曾” as a case study, as shown in Figure 8. This character is structurally complex and shares high visual similarity with characters like “普” and “曹”, making it an ideal candidate to test the model’s discriminative power. In the baseline branch (without semantic matching), the model relies solely on calculating fixed distances in the visual vector space. Since the character “曾” and its specific radical sequence were never seen during training, the global visual embedding fails to capture the internal topological nuances. Consequently, the model essentially performs a nearest-neighbor search based on coarse global shapes, inevitably retrieving morphologically similar but incorrect characters like “普” or “曹”, failing to distinguish the target.
In contrast, our proposed approach actively integrates the extracted auxiliary information with image features, injecting the specific radical sequence of “曾” as a prior constraint. However, the reasoning process reveals a nuanced behavior: relying on a single nearest neighbor (Top-1) is often insufficient for perfect reconstruction. As illustrated in the visualization, the immediate Top-1 retrieval result may only capture partial structural features—for instance, retrieving a character that shares only the top radical or the bottom radical, but fails to embody the complete structure of “曾”. This observation aligns with our hypothesis that zero-shot recognition in a high-dimensional space is rarely a point-to-point match but rather a subspace alignment problem.
The true robustness of our framework emerges from the Top-K Soft Aggregation mechanism. By expanding the retrieval scope to the Top-5 candidates, the model effectively aggregates complementary structural cues from a semantic neighborhood. As shown in the heatmap visualization, the retrieved neighbors (including “普”, “曹”, and “半”) collectively contain all the constituent radicals required to construct “曾” (highlighted in red in Figure 8). While “普” provides the top structure and “半” contributes the middle components, the aggregation module effectively assembles these dispersed semantic fragments. By integrating the intersection of radical information from these Top-$K$ neighbors, the model successfully reconstructs the full semantic identity of “曾”. This case study empirically validates that our method performs structural reasoning rather than simple pattern matching, utilizing the consensus of local radicals to hallucinate the correct prototype even for unseen classes.
Figure 8: Visualization of the Cross-Modal Reasoning Process. We compare the inference outcomes for the unseen character “曾”. The baseline fails due to visual ambiguity with shape-similar characters. Our method leverages radical priors. Notably, while a single neighbor may only match partial radicals, the Top-5 neighbors collectively provide the complete set of necessary radicals (highlighted in red), allowing the Soft Aggregation module to correctly reconstruct the target identity.
@@ 5 Conclusion
In this paper, we presented the Entropy-Aware Structural Alignment Network, a novel framework designed to bridge the gap between seen and unseen classes in zero-shot and few-shot Handwritten Chinese Character Recognition. Addressing the limitations of flat sequence modeling and rigid template matching, we introduced three key innovations: an information-theoretic entropy modulation strategy, a dual-view multi-granularity structural encoding approach, and a Top-K semantic feature fusion mechanism.
Our extensive experiments yield several critical insights: (1) Entropy as a Multiplicative Modulator: We demonstrated that the informational value of radicals follows a heavy-tailed distribution. By modulating positional embeddings via multiplicative interaction based on information entropy, the model successfully prioritizes discriminative roots over ubiquitous particles. This approach proves superior to simple additive embeddings in suppressing high-frequency redundant signals. (2) Adaptive Structural Fusion: The ablation studies confirm that extracting five distinct multi-granularity features (from radical codes to global/local tree views) provides a comprehensive structural description. Furthermore, the adaptive Sigmoid-based GateFusion mechanism effectively balances these heterogeneous cues, allowing the model to dynamically shift focus between global topology and local details based on character complexity. (3) Robustness via Feature Fusion: The visualization of the reasoning process reveals that our Top-K Semantic Feature Fusion strategy effectively mitigates the brittleness of Top-1 retrieval. By injecting the centroid of the semantic neighborhood into the decoder’s query vector, the model achieves a ”structural error correction”, reconstructing the correct identity even when the immediate visual-semantic alignment is ambiguous.
Future Work: While our framework proves effective for isolated character recognition, applying it to end-to-end text line recognition remains a challenging frontier, primarily due to the complexity of character segmentation in cursive streams. Additionally, exploring the applicability of the Entropy-Aware mechanism to other ideographic languages or fine-grained object recognition tasks represents a promising direction for future research.

### v2

@@ 4.1.2 Implementation Details
Our framework is implemented in PyTorch on four NVIDIA RTX 4090 GPUs. The visual stream employs a randomly initialized ResNet-34 backbone, which is trained from scratch on the training set to learn domain-specific handwritten features. The semantic radical encoder is fine-tuned using weights from the CLIP (ViT-B/32) image encoder. This fine-tuning process is conducted prior to the main backbone training, following the protocol detailed in our preliminary work [19]. Consequently, all radical features are extracted via this pre-trained CLIP encoder to construct a static feature dictionary. During the main network training, the CLIP parameters are frozen, serving solely for efficient vector retrieval. The entire model is optimized end-to-end using the Adadelta optimizer with a batch size of 32. We employ a MultiStepLR scheduler, where the learning rate is initialized to 0.1 and decayed to 0.01 and 0.001 at the 15th and 20th epochs, respectively. Regarding key hyperparameters, the neighbor count $K$ for the Top-K Semantic Feature Fusion strategy is set to 5 based on our ablation studies.
@@ 4.2.1 Performance on Zero-shot Character Recognition
The core contribution of this work is the ability to recognize unseen characters by leveraging radical-based semantic priors. We evaluate our method against a comprehensive list of SOTA approaches. Table 1 summarizes the Top-1 accuracy on the ICDAR 2013 target set (Unseen) under the zero-shot setting, where the number of seen training classes $m$ varies from 500 to 2,755. Our Entropy-Aware Structural Alignment Network establishes a new SOTA, consistently outperforming all existing methods across different data regimes.
@@
Superiority over Structural and Decomposition Models: Our method demonstrates a substantial lead over traditional decomposition approaches (e.g., +34.47% over HDE and +25.53% over ACPM), proving that static radical mapping is insufficient for handwritten variance. Furthermore, compared to the latest structural models from 2024 to 2025, our approach significantly outperforms the stroke-based RSST by +20.52% and the hierarchical HierCode by +11.75%. These results highlight a critical insight: while strokes and standard trees capture topology, they often introduce noise. Our Information Entropy Prior effectively suppresses such ubiquitous structural noise, allowing the model to focus on discriminative semantic roots, which proves more robust than stroke-level or unweighted hierarchical features.
Advancement over CLIP-Enhanced and Preliminary Works: A critical comparison is drawn against CLIP-based baselines to validate our specific contributions. Our method surpasses the strong CCR-CLIP by +5.37%, confirming that explicitly modeling the parent-child topology is superior to flat visual-semantic alignment. Most importantly, compared to our preliminary conference version RSE-CLIP (59.80%), which shares the same structural backbone but lacks the entropy mechanism, our proposed framework achieves a remarkable +8.16% improvement. This significant performance leap serves as a direct validation, confirming that the introduction of Entropy-Aware Saliency and Adaptive Top-K Fusion is the decisive factor in resolving the ambiguity of unseen characters.
@@ 4.2.2 Generalization to Full-Set Recognition
While our primary focus is zero-shot learning, we also evaluate the proposed framework on the standard fully supervised setting to verify its general applicability. Table 2 compares our method with SOTA fully supervised approaches based on their auxiliary supervision (e.g., Glyphs, Radicals) and similarity metrics. As shown, our method achieves a highly competitive accuracy of 97.30%, demonstrating that the proposed entropy-aware structural alignment remains robust and effective even when abundant training data is available.
@@
The results also highlight the importance of the matching mechanism. Most existing methods (e.g., HDE, SLD) employ Cosine Similarity, which is computationally efficient but often fails to capture complex non-linear cross-modal dependencies. By introducing Cross-Modal Attention followed by Top-K Aggregation, our framework can better align visual features with semantic embeddings, leading to more robust recognition than fixed-distance metrics.
@@ 4.3 Ablation Studies
4.3.1 Effectiveness of EAPE
To isolate the contribution of the proposed EAPE, we conduct a progressive ablation study comparing it against standard positional encoding schemes on the zero-shot radical recognition task. The baseline employs a simple one-hot radical encoding without any structural position information. We then incrementally incorporate the radical tree topology and various position embedding strategies to measure their specific impacts.
The quantitative results are summarized in Table 4. As observed, the introduction of the Radical Tree topology alone (without explicit position indices) triggers the most significant performance leap, boosting the accuracy from 16.53% to 23.23% at $m=500$. This confirms that the hierarchical connection itself provides essential structural context that a flat baseline lacks. However, simply adding standard geometric encodings—such as Sinusoidal PE or RoPE—to this tree structure yields only marginal gains (e.g., Sinusoidal PE improves by merely $\sim$0.4% over the tree-only baseline). This saturation suggests that treating character structure purely as a geometric sequence is suboptimal, as it fails to account for the varying semantic importance of different nodes. In contrast, our Entropy-Aware mechanism effectively breaks this bottleneck, achieving the highest accuracy of 24.54% ($m=500$) and 55.04% ($m=1500$). By utilizing multiplicative interaction to modulate the positional embeddings, our method amplifies the feature magnitude of high-entropy radicals while dampening ubiquitous particles.
@@ 4.4 Analysis of Few-Shot Adaptability and Data Efficiency
While zero-shot recognition demonstrates the model’s ability to reason from semantic definitions, real-world applications often permit a ”cold start” with a minimal number of support samples. To evaluate the data efficiency of our framework, we investigate the performance gains when providing $N_{s}$ reference samples per unseen category, specifically setting $N_{s}\in\{1,2,5\}$. The performance trends and detailed numerical results are presented in Figure 7.
Rapid Adaptation Performance: Notably, a single visual prototype ($N_{s}=1$) triggers a massive leap: for the $m=500$ split, accuracy surges from 24.54% (Zero-shot) to 92.41%, effectively bridging the majority of the domain gap. Increasing to $N_{s}=2$ further refines accuracy to 96.62%, while $N_{s}=5$ shows clear saturation at 97.93%. This steep ascent followed by diminishing returns confirms our model’s high sample efficiency, achieving robust recognition with as few as two examples. Mechanism of Efficiency: We attribute this capability to the Entropy-Aware Structural Prior. Unlike conventional few-shot learning, which requires multiple samples to statistically infer discriminative features (”Statistical Exploration”), our framework operates with a pre-computed ”structural blueprint.” The information entropy explicitly highlights high-value radicals beforehand, allowing the attention mechanism to immediately ”lock onto” critical regions upon seeing the first sample ($N_{s}=1$). This effectively transforms the learning process into direct ”Structural Registration,” enabling precise alignment with minimal visual evidence.
@@ 4.5 Visualization of Cross-Modal Reasoning Process
To demystify the cross-modal reasoning, we visualize the inference of a challenging unseen character “曾” in Figure 8. The baseline, relying solely on global visual features, fails to capture internal topology and misclassifies it as morphologically similar characters (e.g., “普” or “曹”). In contrast, our Entropy-Aware network imposes radical-level constraints. Crucially, the visualization reveals that while a single Top-1 retrieval may only match partial components, the Top-5 semantic neighbors (including “普”, “曹”, and “半”) collectively contain all the constituent radicals required to construct “曾” (highlighted in red). The Soft Aggregation mechanism effectively assembles these dispersed semantic fragments, successfully reconstructing the full target identity from the consensus of local structural cues.
Figure 8: Visualization of the reasoning process for the unseen character “曾”. The baseline fails due to visual ambiguity with shape-similar characters. Our method leverages radical priors. Notably, while a single neighbor may only match partial radicals, the Top-5 neighbors collectively provide the complete set of necessary radicals (highlighted in red), allowing the Soft Aggregation module to correctly reconstruct the target identity.
@@ 5 Conclusion
In this paper, we presented the Entropy-Aware Structural Alignment Network, a novel framework designed to bridge the gap between seen and unseen classes in zero-shot and few-shot Handwritten Chinese Character Recognition. Addressing the limitations of flat sequence modeling and rigid template matching, we introduced three key innovations: an information-theoretic entropy modulation strategy, a dual-view multi-granularity structural encoding approach, and a Top-K semantic feature fusion mechanism. Our extensive experiments yield several critical insights: (1) Entropy as a Multiplicative Modulator: We demonstrated that the informational value of radicals follows a heavy-tailed distribution. By modulating positional embeddings via multiplicative interaction based on information entropy, the model successfully prioritizes discriminative roots over ubiquitous particles. This approach proves superior to simple additive embeddings in suppressing high-frequency redundant signals. (2) Adaptive Structural Fusion: The ablation studies confirm that extracting five distinct multi-granularity features (from radical codes to global/local tree views) provides a comprehensive structural description. Furthermore, the adaptive Sigmoid-based GateFusion mechanism effectively balances these heterogeneous cues, allowing the model to dynamically shift focus between global topology and local details based on character complexity. (3) Robustness via Feature Fusion: The visualization of the reasoning process reveals that our Top-K Semantic Feature Fusion strategy effectively mitigates the brittleness of Top-1 retrieval. By injecting the centroid of the semantic neighborhood into the decoder’s query vector, the model achieves a ”structural error correction”, reconstructing the correct identity even when the immediate visual-semantic alignment is ambiguous.
Future Work: While our framework proves effective for isolated character recognition, applying it to end-to-end text line recognition remains a challenging frontier, primarily due to the complexity of character segmentation in cursive streams. Additionally, exploring the applicability of the Entropy-Aware mechanism to other ideographic languages or fine-grained object recognition tasks represent a promising direction for future research.
6 CRediT authorship contribution statement
Qiuming Luo: Conceptualization, Methodology, Resources, Supervision. Tao Zeng: Investigation, Software, Formal analysis. Feng Li: Data Curation, Funding acquisition. Heming Liu: Visualization, Funding acquisition. Rui Mao: Validation, Writing - Review & Editing. Chang Kong: Writing - Original Draft, Project administration, Supervision.
7 Data availability
This work used publicly datasets.
8 Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.